K is the midpoint of line segment JL. J is located at (7,2) and L is located at (4,4). What are the coordinates of K?
Answer by stanbon(75887)

K is the midpoint of line segment JL. J is located at (7,2) and L is located at (4,4). What are the coordinates of K?
---------------
Let the coordinates be (x,y)
------
Equations:
x = (7+4)/2 = 5.5
y = (4+2)/2 = 3
=====================
Cheers,
Stan H.
=====================